# Unified Microphone Conversion: Many-to-Many Device Mapping via Feature-wise Linear Modulation

### v2

@@ 1 Introduction
Motivated by this limitation, we propose a many-to-many device mapping approach utilizing a CycleGAN framework combined with Feature-wise Linear Modulation (FiLM) [6], incorporating frequency response data of recording devices. We hypothesize that integrating device frequency response information via FiLM can accurately specify the desired inter-domain mappings while maintaining consistent acoustic information. Specifically, we modulate the channel-wise statistics of the generator’s intermediate embeddings by incorporating frequency-response-related embeddings, allowing the model to capture device-specific transformations. By not restricting the model to particular source–target device pairs, this method provides a scalable and adaptable solution to address device variability in sound event classification. Lastly, we introduce a synthetic frequency difference generation algorithm to enhance the scalability of the Unified Microphone Conversion network, enabling more efficient training of SEC models across diverse domains without needing additional frequency response data.
@@ 2 Related Work
Recently, a Microphone Conversion approach [11] utilized the CycleGAN framework to address device variability by generating synthetic spectrograms that mimic recordings from various devices. While it outperformed all previous methods and established the state-of-the-art for handling device variability across multiple devices, its reliance on one-to-one domain mappings limits scalability, as a separate model must be trained for each device pair—making it a strong but impractical baseline for comparison. This requirement becomes increasingly challenging with the expanding array of recording devices, especially if new data becomes available or additional devices are introduced, as it necessitates retraining the network. Moreover, since the DCASE dataset lacks device impulse response information, we conduct our experiments on a more suitable dataset containing real device IRs and greater class and device diversity.
@@ 3.1 Unified Microphone Conversion
Our methodology refines the CycleGAN framework for the task of simulating device variability in audio recordings. In contrast to the conventional CycleGAN setup, which employs two generators and two discriminators, our modified architecture consolidates two generators into a single unified generator augmented by outputs of a FiLM encoder. This generator is supported by $k$ discriminators, each assigned to one of $k$ distinct device domains. We adopt the architecture for the generator and discriminator from Ryu et al. [11].
@@ 4.1 Dataset
We extend the dataset by adding frequency response data of each device, defined as log Mel spectra (128 Mel bands) from 200-ms segments of recorded impulses. The Kronecker delta function is used to generate the impulse with readily available mobile phones, providing a cost-effective alternative to more rigorous methods for characterizing the frequency response.

### v3

@@ Abstract
footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Motivated by this limitation, we propose a many-to-many device mapping approach utilizing a CycleGAN framework combined with Feature-wise Linear Modulation (FiLM) [6], incorporating frequency response data of recording devices. We hypothesize that integrating device frequency response information via FiLM can accurately specify the desired inter-domain mappings while maintaining consistent acoustic information. Specifically, we modulate the channel-wise statistics of the generator’s intermediate embeddings by incorporating frequency-response-related embeddings, allowing the model to capture inter-device transformations. By not restricting the model to particular source–target pairs, this method provides a scalable and adaptable solution to address device variability in sound event classification. Moreover, we introduce a synthetic frequency difference generation algorithm to enhance the scalability of the Unified Microphone Conversion network, enabling more efficient training of SEC models across diverse domains without needing additional frequency response data.
@@ 2 Related Work
Recently, a Microphone Conversion approach [11] utilized the CycleGAN framework to address device variability by generating synthetic spectrograms that mimic recordings from various devices. While it outperformed all previous methods and established the state-of-the-art for handling device variability across multiple devices, its reliance on one-to-one domain mappings limits scalability, as a separate model must be trained for each device pair—making it a strong but impractical baseline for comparison. This requirement becomes increasingly challenging with the expanding array of recording devices, especially if new data becomes available or additional devices are introduced, as it necessitates retraining the network.
@@ 3.1 Unified Microphone Conversion
In contrast to the conventional CycleGAN setup, which employs two generators and two discriminators, our modified architecture consolidates two generators into a single unified generator augmented by outputs of a FiLM encoder. This generator is supported by $k$ discriminators, each assigned to one of $k$ distinct device domains. We adopt the architectures for the generator and discriminator from Ryu et al. [11].
@@ 4.1 Dataset
We extend the dataset by adding frequency response data of each device, defined as log Mel spectra (128 Mel bands) from 200-ms segments of recorded impulses. The Kronecker delta function is used to generate the impulse with readily available mobile phones, providing a cost-effective alternative to more rigorous methods for characterizing the frequency response. Moreover, as the DCASE dataset lacks impulse response information, our dataset-which includes real device IRs and greater class and device diversity-is more suitable for our experiments.